### v1

@@ 4. Nice Words
In this section, we prove Lemma 3.4. To this effect, given $p\geq 1$, let $K_{p}\subset({\mathbb{Z}}^{m})^{(p+1)}$ be the set of all indices $\bf{k_{1},\ldots,k_{p}}$ such that $\ldots\bf{[k_{0}V,k_{1}V],\ldots],k_{p}V]}$ can be written of the form $\log(W_{1}W_{2}^{-1})\text{ mod }\mathfrak{g}^{(p+1)}$ for two words $W_{1}$ and $W_{2}$ of equal length in $\langle F\rangle$.
@@
Given an $m$-tuple $F$, choose elements $g_{1},g_{2}\in\langle F\rangle$ and we define new elements $\hat{g}^{(p)}_{L}$ and $\hat{g}^{(p)}_{R}$ inductively as follows. Let $w_{i},w_{i}^{\prime}\in\langle F\rangle$ for each $i\in{\mathbb{N}}_{0}$
(4.1) $\hat{g}_{L}^{(0)}:=w_{0},\quad\hat{g}_{R}^{(0)}:=w_{0}^{\prime},$
(4.2) $\hat{g}_{L}^{(p)}:=\hat{g}_{L}^{(p-1)}w_{p}\hat{g}_{R}^{(p-1)},\quad\hat{g}_{R}^{(p)}:=\hat{g}_{R}^{(p-1)}w_{p}\hat{g}_{L}^{(p-1)}.$
As an initial case, $g_{L}^{(0)},g_{R}^{(0)}]=g_{L}^{(1)}(g_{R}^{(1)})^{-1}.$ By direct computation, we obtain
(4.3) $\begin{split}\hat{g}_{L}^{(p)}(\hat{g}_{R}^{(p)})^{-1}&=\hat{g}_{L}^{(p-1)}w_{p}\hat{g}_{R}^{(p-1)}(\hat{g}_{R}^{(p-1)}w_{p}\hat{g}_{L}^{(p-1)})^{-1}\\ &=[\hat{g}_{L}^{(p-1)}(\hat{g}_{R}^{(p-1)})^{-1},\hat{g}_{R}^{(p-1)}w_{p}]\\ &=[\ldots[[\hat{g}_{L}^{(1)}(\hat{g}_{R}^{(1)})^{-1},\hat{g}_{R}^{(1)}w_{2}],\hat{g}_{R}^{(2)}w_{3}],\ldots],\hat{g}_{R}^{(p-1)}w_{p}]\\ &=[\ldots[[[\hat{g}_{L}^{(0)},\hat{g}_{R}^{(0)}],\hat{g}_{R}^{(1)}w_{2}],\ldots],\hat{g}_{R}^{(p-1)}w_{p}].\end{split}$
By the BCH formula, we write
(4.4) $\log\hat{g}_{L}^{(p)}(\hat{g}_{R}^{(p)})^{-1}=[\ldots[[\log[\hat{g}_{L}^{(0)},\hat{g}_{R}^{(0)}],\log(\hat{g}_{R}^{(1)}w_{2})],\ldots],\log(\hat{g}_{R}^{(p-1)}w_{p})]\text{ mod }\mathfrak{g}^{(p+1)}.$
Letting $\bf{k_{i}}$ be the tuple storing in position $j$ the number of times $g_{j}$ appears in $g_{R}^{(i-1)}w_{i}$, we get that this reduces to
(4.5) $\log\hat{g}_{L}^{(p)}(\hat{g}_{R}^{(p)})^{-1}=[\ldots[\mathbf{k_{0}V,k_{1}V],\ldots,k_{p}V]}\text{ mod }\mathfrak{g}^{(p+1)}.$
Now we wish to understand which sequences $\bf{k_{0},\ldots,k_{p}}$ are possible to choose. Note that $\bf{k_{0}}$ and $\bf{k_{1}}$ are completely free to choose. Meanwhile, for $i\geq 2$, we have $\bf{k_{i}\geq k_{i-1}}$ as the only restraint arising from this construction. It is clear that the set of all tuples satisfying this restraint when interpreted as a subset of ${\mathbb{R}}^{mp}$ is Zariski dense. It follows that $K_{p}$ is a Zariski dense set. ∎
Proof of Lemma 3.4.
We claim that there are words $W_{1}$ and $W_{2}$ such that the coordinates of $\pi_{p}(\log(W_{1}W_{2}^{-1})$ are linearly independent polynomials in $\alpha$. Assume by way of contradiction that for all words $W_{1},W_{2}$, these polynomials are linearly dependent. In this case $K_{p}\subset\mathcal{D}(\mathcal{H}_{m,p}).$ By Lemma 4.1, $\mathcal{H}_{m,p}$ is degenerate contradicting the assumption that $\mathfrak{g}$ is mm–great. The claim follows.
Next, by Theorem 2.5, since the coordinates of $\pi_{p}(W_{1}W_{2}^{-1})$ are linearly independent, for almost every value of the $g_{i}$, $\pi_{p}(W_{1}W_{2}^{-1})\in DC(\gamma,\tau).$ This implies the Lemma. ∎
@@ 5. Great Groups
Now that we have established that $m$-greatness implies rapid mixing of almost all random walks on $m$ elements, we wish to show that this condition is not overly restrictive. To that effect, we show that many special Lie algebras are 2-great, and that every step-$s$ Lie algebra is $s$-great. This will prove Theorems 1.2 and 1.3. Finally, we present an example of a group that is not 2-great.
@@
Here we prove that all quasi-abelian, triangular and step-3 or lower Lie algebras are 2-great. Our approach in the first two cases will be to choose a value of $k$ at which $\mathcal{H}_{2,p}(k,\alpha)$ has linearly independent polynomials. We shall use $\bar{k}_{p}=((1,0),(0,1)^{p})$. We will then choose a particular monomial and, by ignoring all variables not appearing in that monomial, find that monomial appears in exactly one coordinate. This will be sufficient to establish linear independence.
@@
To show this, we must show that $\mathcal{H}_{2,p}$ is nondegenerate for all $1\leq p\leq s$. To that effect, we will let
@@
and compute the coefficient of $\alpha_{0,i}\beta_{0}^{p}$ in $\mathcal{H}_{2,p}(\bar{k},\alpha).$ In particular, we see that if $(i,p)\in J$ then this coefficient is equal to $Y_{i,p}$. This implies that the polynomials of $\mathcal{H}_{2,p}(\bar{k},\alpha)$ are linearly independent since each one includes a unique monomial. Thus $\mathcal{H}_{2,p}$ is not degenerate for any $p$, and hence any quasi-abelian Lie algebra is 2-great. ∎
@@
Let $V_{1}=\sum_{i=1}^{n-1}\alpha_{0i}E_{i,i+1}\quad\mathrm{mod}\quad\mathfrak{g}^{(1)}$ and $V_{2}=\sum_{i=1}^{n-1}\beta_{0i}E_{i,i+1}\quad\mathrm{mod}\quad\mathfrak{g}^{(1)}.$ We compute the coefficient of $\alpha_{i}\beta_{i+1},\ldots\beta_{i+p}$ in $\mathcal{H}_{2,p}(\bar{k}_{p},\alpha)$ for $1\leq i\leq s-p$ and $1\leq p\leq s$ to be $E_{p,i}$. This implies that the polynomials of $\mathcal{H}_{2,p}(\bar{k},\alpha)$ are linearly independent since each one includes a unique monomial. Thus, we see that $\mathcal{H}_{2,p}$ is not degenerate for any $p$, and conclude that $\mathfrak{t}_{s}$ is 2-great. ∎
@@
Now, we show that step-3 Lie algebras are also necessarily 2-great. This is the best we can do in terms of step as there exist step-4 Lie algebras that are not 2-great as we shall see.
@@
Since terms of the form $\alpha_{i_{1}}^{2}\beta_{i_{2}}$ occur with coefficient $X_{i_{1}},X_{i_{2}}],X_{i_{1}}]$, we get
@@
Since the matrix defining (5.3)–(5.4) is non degenerate we conclude that
@@
Combining Propositions 5.1, 5.2, and 5.5 with Theorem 3.7 gives us Theorem 1.3.
@@ 5.2. General Lie algebras
We prove that $\mathcal{H}_{s,s-1}$ is non degenerate. Let $V_{i}=\sum_{j=1}^{n_{0}}\alpha_{ij}X_{j}\text{ mod }\mathfrak{g}^{(1)},$ and consider $\bf{[\ldots[k_{0}V,k_{1}V],\ldots],k_{s-1}V]}$. Let $\bar{k}$ denote the conveniently selected parameters: $\mathbf{k_{0}}=(1,0,\ldots,0),\mathbf{k_{1}}=(0,1,\ldots,0),\ldots,\mathbf{k_{s-1}}=(0,0,\ldots,1)$. We compute that the coefficient of $\alpha_{0i_{0}}\cdots\alpha_{s-1i_{s-1}}$ in $\mathcal{H}_{s,s-1}(\bar{k},\alpha)$ is exactly $\ldots[[X_{i_{0}},X_{i_{1}}],\ldots,X_{i_{s-1}}]$. Since each element of the basis for $\mathfrak{g}^{(s-1)}$ can be written of the form $\ldots[[X_{i_{0}},X_{i_{1}}],\ldots,X_{i_{s-1}}]$, we get that each coordinate polynomial of $\mathcal{H}_{s,s-1}(\bar{k},\alpha)$ has a monomial appear which is unique to that coordinate. Thus, the polynomials at $\bar{k}$ are linearly independent, so $\mathcal{H}_{s,s-1}$ is non-degenerate. Since every $\mathfrak{g}/\mathfrak{g}^{(i+1)}$ is step–$s$ or lower, the same argument will show that $\mathcal{H}_{s,i}$ is non degenerate for all $i<s.$ We conclude that $\mathfrak{g}$ is ss–great. ∎
Combining Proposition 5.5 with Theorem 3.7 gives us Theorem 1.2. We also show that $m$-greatness is preserved by taking products and factors.
@@
For clarity, in this proof we will let $\mathcal{H}^{\mathfrak{g}}_{m,p}$ be the map $\mathcal{H}_{m,p}$ associated to the Lie group $\mathfrak{g}.$ Similarly, let $\alpha^{\mathfrak{g}}$ be the Malcev coordinates in the basis of $\mathfrak{g}$ and $\mathcal{D}^{\mathfrak{g}}$ be the set of $k$ such that $\mathcal{H}^{\mathfrak{g}}_{m,p}$ is linearly dependent.
For the first part, observe that $\mathcal{H}_{m,p}^{\mathfrak{g}_{1}\times\mathfrak{g}_{2}}(k,\alpha^{\mathfrak{g}_{1}\times\mathfrak{g}_{2}})=(\mathcal{H}_{m,p}^{\mathfrak{g}_{1}}(k,\alpha^{\mathfrak{g}_{1}}),\mathcal{H}_{m,p}^{\mathfrak{g}_{2}}(k,\alpha^{\mathfrak{g}_{2}}))$. Since both of the $\mathfrak{g}_{i}$’s are $m$-great, $\mathcal{D}_{\mathfrak{g}_{i}}$ is a positive codimension variety. Thus, there exists a $k$ for which $\mathcal{H}_{m,p}^{\mathfrak{g}_{1}}(k,\alpha^{\mathfrak{g}_{1}})$ and $\mathcal{H}_{m,p}^{\mathfrak{g}_{2}}(k,\alpha^{\mathfrak{g}_{2}})$ both have linearly independent coordinate polynomials. It follows that for this choice of $k$, $\mathcal{H}_{m,p}^{\mathfrak{g}_{1}\times\mathfrak{g}_{2}}(k,\alpha^{\mathfrak{g}_{1}\times\mathfrak{g}_{2}})$ has the same property. Thus $\mathcal{H}_{m,p}^{\mathfrak{g}_{1}\times\mathfrak{g}_{2}}$ is non-degenerate for any $p$, so $\mathfrak{g}_{1}\times\mathfrak{g}_{2}$ is $m$-great.

### v2

@@ 4. Nice Words
In this section, we prove Proposition 3.4. Our approach is to construct words in the semigroup $\langle F\rangle$ that differ by a nonzero element of $G^{(p)}$ for each $p$. In particular, we will find words that differ by a term that can be expressed as a $p$-fold bracket of elements of $\langle F\rangle$. Then, we will observe that a formula for this difference may be given as a polynomial in the exponential coordinates of the $g_{i}$. As long as the polynomials arising in this manner are linearly independent, we can apply Theorem 2.5 to conclude that, for a full measure set of walks, the words we construct act Diophantinely on the appropriate level, and thus we have nice words. This need for linear independence was the direct motivation for Definition 3.2. Once we have linear independence, Theorem 2.5 will guarantee the existence of a full measure of the set of $m$-tuples that act Diophanitnely and thus rapidly mix.
To begin, given $p\geq 1$, let $K_{p}\subset({\mathbb{Z}}^{m})^{(p+1)}$ be the set of all indices $\bf{k_{0},\ldots,k_{p}}$ such that $\ldots\bf{[k_{0}V,k_{1}V],\ldots],k_{p}V]}$ can be written of the form $\log(W_{1}W_{2}^{-1})\text{ mod }\mathfrak{g}^{(p+1)}$ for two words $W_{1}$ and $W_{2}$ of equal length in $\langle F\rangle$. This is the set of $p$-nested commutators that arise as the logarithm of the difference between two elements of our walk at some time. Thus $K_{p}$ can heuristically be seen as the set of elements of $\mathfrak{g}_{p}$ available to us in trying to get a self-cancellation between a character on level $p$ pushed forward by two words in our random walk. In constructing elements of $K_{p}$, the element $k_{ij}$ will count how many times $g_{j}$ appears in the $i^{th}$ step of the construction of the words differing by $\mathcal{H}_{p}(k,\alpha)$. For instance, in the setting of two generators $g_{1}$ and $g_{2}$, since
$\log(g_{1}g_{2}(g_{2}g_{1})^{-1})=\log([g_{1},g_{2}])=[{\bf k_{0}V,k_{1}V}]$
for ${\bf k}_{0}=(1,0)$ and ${\bf k}_{1}=(0,1)$, we have that $((1,0),(0,1))\in K_{2}$. In terms of this set, our goal is to find an intersection between $K_{p}$ and the complement of $\mathcal{D}(\mathcal{H}_{m,p})$ i.e. a difference between two words in our walk with linearly independent polynomials. The following lemmas will demonstrate a method for finding elements of $K_{p}$ that will be rich enough that as long as $\mathcal{D}(\mathcal{H}_{m,p})\neq{\mathbb{R}}^{n_{p}}$, $K_{p}$ will not be contained in $\mathcal{D}(\mathcal{H}_{m,p})$ as we desire.
Lemma 4.1.
For any given $w_{-1},\ldots w_{p}\in\langle\mathcal{F}\rangle$ with $|w_{-1}|=|w_{0}|$, there exist words of equal length $W_{1}$ and $W_{2}$ such that
$W_{2}W_{1}^{-1}\in\mathfrak{g}_{p}.$
Proof.
Our idea is to inductively build words $a_{i}$ and $b_{i}$ in $\langle F\rangle$ such that $h_{i}:=a_{i}b_{i}^{-1}\in G^{(i)}$ and to do so with as much flexibility as possible. To begin with, we set
$a_{0}=w_{-1}\quad b_{0}=w_{0},$
and
$a_{i}=a_{i-1}w_{i}b_{i-1}\quad\text{and}\quad b_{i}=b_{i-1}w_{i}a_{i-1}.$
Since we took $|w_{-1}|=|w_{0}|$, this construction will always give us words $a_{i}$ and $b_{i}$ satisfying $|a_{i}|=|b_{i}|$. Observe that
$h_{1}=a_{0}w_{1}b_{0}a_{0}^{-1}w_{1}^{-1}b_{0}^{-1}=[a_{0}b_{0}^{-1},b_{0}w_{1}],$
and
$h_{i}=a_{i-1}w_{i}b_{i-1}(b_{i-1}w_{i}a_{i-1})^{-1}=a_{i-1}(b_{i-1}^{-1}b_{i-1})w_{i}b_{i-1}a_{i-1}^{-1}w_{i}^{-1}b_{i-1}^{-1}=[a_{i-1}b_{i-1}^{-1},b_{i-1}w_{i}]$
Since $a_{i-1}b_{i-1}^{-1}=h_{i-1}$, we can rewrite this as
$h_{i}=[h_{i-1},b_{i-1}w_{i}].$
Thus, $h_{1}$ can be expressed as a single bracket, and $h_{i}$ can be expressed as a bracket between $h_{i-1}$ and another group element, so inductively $h_{i}$ can always be expressed as an $i$-fold bracket. In particular, we get the formula
(4.1) $h_{p}=[\ldots[[a_{0}b_{0}^{-1},b_{0}w_{1}],b_{1}w_{2}],\ldots,b_{p-1}w_{p}],$
so $h_{p}\in\mathfrak{g}^{(p)}$ as desired. ∎
Lemma 4.2.
For any $k$ satisfying $3k_{ij}<k_{i+1j}$, there exist words of equal length $W_{1},W_{2}$ such that
$\log(W_{2}W_{1}^{-1})=[\ldots\bf{[k_{0}V,k_{1}V],\ldots],k_{p}V]}\mod\mathfrak{g}^{(p+1)}.$
Proof.
@@
By the previous Lemma, for any $w_{-1},\ldots w_{p}$ with $|w_{-1}|=|w_{0}|$, we can build words that differ by $h_{p}\in\mathfrak{g}^{(p)}$. Moreover, $h_{p}$ can be written as $a_{p}b_{p}^{-1}$ for two words of equal length in $\langle F\rangle$. If we can find ${\bf k_{0}},{\bf k_{1}},\ldots,{\bf k_{p}}$ such that
$\log h_{p}=[\ldots\bf{[k_{0}V,k_{1}V],\ldots],k_{p}V]},$
then $({\bf k_{0}},{\bf k_{1}},\ldots,{\bf k_{p}})\in K_{p}$. Motivated by this, we take the logarithm of both sides of (4.1) to get
$\log(h_{p})=\log([\ldots[[a_{0}b_{0}^{-1},b_{0}w_{1}],b_{1}w_{2}],\ldots,b_{p-1}w_{p}]).$
The BCH formula allows us to move the logarithm within the brackets modulo some error that belongs to $\mathfrak{g}^{(p+1)}$, so
(4.2) $\log(h_{p})=[\ldots[[\log(a_{0}b_{0}^{-1}),\log(b_{0}w_{1})],\log(b_{1}w_{2})],\ldots,\log(b_{p-1}w_{p})]\mod\mathfrak{g}^{(p+1)}.$
Now, define a function ${\bf c}:\langle F,F^{-1}\rangle\rightarrow{\mathbb{Z}}^{m}$ such that ${\bf c}_{i}(W)$ is the number of times $g_{i}$ appears in the word $W$ with inverses counted as negatives. Define ${\bf k_{0}}={\bf c}(a_{0}b_{0}^{-1})$ and ${\bf k_{i}}={\bf c}(b_{i-1}w_{i})$. Applying BCH again, we have that $\log(b_{i-1}w_{i})=\sum_{i=1}^{m}{\bf k_{i}}V_{i}\mod\mathfrak{g}^{(1)}$. Applying this fact to (4.2) and using the fact that the $p$-fold bracket of a $\mathfrak{g}^{(1)}$ error is in $\mathfrak{g}^{p+1}$, we get
$\log(h)=[\ldots[\sum_{i=1}^{m}{\bf k_{0i}}V_{i},\sum_{i=1}^{m}{\bf k_{1i}}V_{i}],\ldots,\sum_{i=1}^{m}{\bf k_{pi}}V_{i}]=[\ldots\bf{[k_{0}V,k_{1}V],\ldots],k_{p}V]}\mod\mathfrak{g}^{(p+1)}.$
Now, we have ${(\bf k_{0},\ldots,k_{p}})\in K_{p}$ which depends on our initial choice of $w_{-1},\ldots,w_{p}$. We wish to understand which sequences $\bf{k_{0},\ldots,k_{p}}$ are possible to get as a result of that choice. Note that $\bf{k_{0}}$ and $\bf{k_{1}}$ can be taken to be any vector of nonnegative integers simply by making an appropriate choice of $w_{-1},w_{0},$ and $w_{1}$. Meanwhile, for $i\geq 2$, we have ${\bf k_{i}}={\bf c}(b_{i-1})+{\bf c}(w_{i})$. Since by our choice of $w_{i}$ we can make ${\bf c}(w_{i})$ any nonnegative integer vector we like, this means we can choose any $\bf k_{i}$, subject to ${\bf k_{i}}\geq{\bf c}(b_{i-1})$. Since ${\bf c}(b_{i-1})=2{\bf c}(b_{i-2})+{\bf c}(w_{i-1})\leq 3{\bf k}_{i-1}$, if ${\bf k_{i}}\geq{\bf k_{i-1}}$, then ${(\bf k_{0},\ldots,k_{p}})\in K_{p}$. ∎
@@
By the previous lemma, $\{k\in({\mathbb{Z}}^{m})^{p+1}:3k_{ij}<k_{i+1j}\}$ is contained in $K_{p}$. It is clear that this set, when interpreted as a subset of ${\mathbb{R}}^{mp}$, is Zariski dense. It follows that $K_{p}$ is a Zariski dense set. ∎
Proof of Proposition 3.4.
We claim that there are words $W_{1}$ and $W_{2}$ of equal length such that the coordinates of $\pi_{p}(\log(W_{1}W_{2}^{-1}))$ are linearly independent polynomials in $\alpha$. Assume by contradiction that for all words $W_{1},W_{2}$, these polynomials are linearly dependent. That implies that any words that differ by $p$-nested commutators have those commutators’ polynomials linearly dependent. Since $K_{p}$ consists of the set of such realized commutators, this means that $K_{p}\subset\mathcal{D}(\mathcal{H}_{m,p}).$ Since $\mathcal{D}(\mathcal{H}_{m,p})$ is an algebraic variety and by Lemma 4.3 $K_{p}$ is Zariski dense, this implies that $\mathcal{D}(\mathcal{H}_{m,p})={\mathbb{R}}^{n_{p}}$. Thus, $\mathcal{H}_{m,p}$ is degenerate, contradicting the assumption that $\mathfrak{g}$ is $m$-great. The claim follows.
Next, by Theorem 2.5, since the coordinate polynomials of $\pi_{p}(W_{1}W_{2}^{-1})$ are linearly independent, for almost every value of the $\alpha_{ij}$, $\pi_{p}(W_{1}W_{2}^{-1})\in DC(\gamma,\tau).$ That means that for almost every $m$-tuple of $g_{i}$’s $\pi_{p}(W_{1}W_{2}^{-1})\in DC(\gamma,\tau).$ This implies the Proposition. ∎
@@ 5. Great Groups
Now that we have established that $m$-greatness implies rapid mixing of almost all random walks on $m$ elements, we wish to show that this condition is not overly restrictive. To that effect, we show that many special Lie algebras are 2-great, and that every step-$s$ Lie algebra is $s$-great. This will prove Theorems A and B. Finally, we present an example of a group that is not 2-great.
We now briefly discuss how we will approach the task of proving that a Lie algebra is $m$-great. Given a Lie algebra $\mathfrak{g}$, we will pick some concrete $k_{p}$ for which we will show that $\mathcal{H}_{m,p}(k_{p},\alpha)$ has linearly independent polynomials i.e. $k_{p}\notin\mathcal{D}(\mathcal{H}_{m,p})$. To show this linear independence, we let $\{P_{1},\ldots,P_{n_{p}}\}$ be the set of $n_{p}$ coordinate polynomials associated to the $X_{i}^{(p)}$. If there is a monomial $M_{i}=\alpha_{i_{0}j_{0}}\ldots\alpha_{i_{p}j_{p}}$ with nonzero coefficient in $P_{i}$ but zero coefficient in each $P_{j}$ when $j\neq i$, we call $M_{i}$ a unique monomial of $P_{i}$. If each $P_{i}$ has a unique monomial, this is sufficient to conclude that the $P_{i}$ form a linearly independent set. By finding a unique monomial of each of the $P_{i}$, we can show that the $P_{i}$ form a linearly independent set. Thus, $\bar{k}_{p}\notin\mathcal{D}(\mathcal{H}_{2,p})$ and $\mathcal{H}_{2,p}$ is non-degenerate. From there, we will conclude 2-greatness.
@@
Here we prove that all quasi-abelian, triangular and step-3 or lower Lie algebras are 2-great. Throughout this section, we shall denote $\bar{k}_{p}=((1,0),(0,1)^{p})$.
@@
To show this, we must show that $\mathcal{H}_{2,p}$ is nondegenerate for all $0\leq p\leq s-1$. We recall that we defined a basis $\{X,Y_{i,j}\}_{(i,j)\in J}$ in Section 2.4. We will let
@@
and let $P_{1},\ldots,P_{n_{p}}$ be the coordinate polynomials in $\alpha$ of $\mathcal{H}_{2,p}(\bar{k}_{p},\alpha)$. We will show that if $P_{i}$ is the polynomial corresponding to the $Y_{i,p}$ coordinate then $\alpha_{0,i}\beta_{0}^{p}$ is a unique monomial of $P_{i}$. This is because the coefficient of $\alpha_{0,i}\beta_{0}^{p}$ in $\ldots[V_{1},V_{2}],\ldots V_{2}]$ is exactly $\ldots[[Y_{0,i},X],X],\ldots,X]=Y_{i,p}$. This means that each $P_{i}$ has a unique monomial, so they form a linearly independent set of polynomials. Thus $\bar{k}\notin\mathcal{D}(\mathcal{H}_{2,p})$, so $\mathcal{H}_{2,p}$ is non-degenerate. Since $p$ is arbitrary, any quasi-abelian Lie algebra is 2-great. ∎
@@
We recall that we defined a basis $\{E_{ij}\}$ in Section 2.4. Let
$V_{1}=\sum_{i=1}^{n-1}\alpha_{0i}E_{i,i+1}\quad\mathrm{mod}\quad\mathfrak{g}^{(1)}\quad\text{ and }\quad\displaystyle V_{2}=\sum_{i=1}^{n-1}\beta_{0i}E_{i,i+1}\quad\mathrm{mod}\quad\mathfrak{g}^{(1)}.$
Let $P_{1},\ldots,P_{n_{p}}$ be the coordinate polynomials in $\alpha$ of $\mathcal{H}_{2,p}(\bar{k}_{p},\alpha)$. We will show that if $P_{i}$ is the polynomial corresponding to the $E_{i,i+p+1}$ coordinate, then $\alpha_{i}\beta_{i+1},\ldots\beta_{i+p}$ is a unique monomial of $P_{i}$. This is because the coefficient of $\alpha_{i}\beta_{i+1},\ldots\beta_{i+p}$ in $\ldots[V_{1},V_{2}],\ldots V_{2}]$ is exactly $\ldots[E_{i,i+1},E_{i+1,i+2}],\ldots,E_{i+p,i+p+1}]=E_{i,i+p+1}$. This means that each $P_{i}$ has a unique monomial, so they form a linearly independent set of polynomials. Thus, $\bar{k}\notin\mathcal{D}(\mathcal{H}_{2,p})$ and so $\mathcal{H}_{2,p}$ is non-degenerate. Since $p$ is arbitrary, any triangular Lie algebra is 2-great. ∎
@@
Now, we show that step-3 Lie algebras are also necessarily 2-great. This is the best we can do in terms of step as there exist step-4 Lie algebras that are not 2-great as we shall see. Our proof technique will be similar to the proofs of $2$-greatness that we have already presented in that we begin by fixing some choice of $k$ which we show to be in the complement of $\mathcal{D}(\mathcal{H}_{2,2})$. It will differ in that we will show that the existence of a linear dependence between the polynomials of $\mathcal{H}_{2,2}$ contradicts the fact that $\mathfrak{g}$ is step 3 rather than directly showing linear independence by identifying unique monomials of each polynomial.
@@
From this we deduce that if a given monomial $\alpha_{i_{1}}\beta_{i_{2}}\alpha_{i_{3}}$ has coefficient $W_{i_{1}i_{2}i_{3}}\in\mathfrak{g}$ in the expression $V_{1},V_{2}],V_{1}]$, then $W_{i_{1}i_{2}i_{3}}\in Ker(\lambda)$. Since terms of the form $\alpha_{i_{1}}^{2}\beta_{i_{2}}$ occur with coefficient $X_{i_{1}},X_{i_{2}}],X_{i_{1}}]$, we get
@@
Since the matrix defining (5.3)–(5.4) is non-degenerate, we conclude that
@@
Combining Propositions 5.1, 5.2, and 5.5 with Theorem 3.7 gives us Theorem B.
@@ 5.2. General Lie algebras
We prove that $\mathcal{H}_{s,s-1}$ is non-degenerate. Let $V_{i}=\sum_{j=1}^{n_{0}}\alpha_{ij}X_{j}\text{ mod }\mathfrak{g}^{(1)},$ and consider $\bf{[\ldots[k_{0}V,k_{1}V],\ldots],k_{s-1}V]}$. Let $\bar{k}$ denote the conveniently selected parameters: $\mathbf{k_{0}}=(1,0,\ldots,0),\ \mathbf{k_{1}}=(0,1,\ldots,0),\ldots,\mathbf{k_{s-1}}=(0,0,\ldots,1)$. We compute that the coefficient of $\alpha_{0i_{0}}\cdots\alpha_{s-1i_{s-1}}$ in $\mathcal{H}_{s,s-1}(\bar{k},\alpha)$ is exactly $\ldots[[X_{i_{0}},X_{i_{1}}],\ldots,X_{i_{s-1}}]$, so $\alpha_{0i_{0}}\cdots\alpha_{s-1i_{s-1}}$ will be a unique monomial for the corresponding coordinate polynomial of $\mathcal{H}_{s,s-1}$. Since each element of the basis for $\mathfrak{g}^{(s-1)}$ can be written of the form $\ldots[[X_{i_{0}},X_{i_{1}}],\ldots,X_{i_{s-1}}]$, we get that each coordinate polynomial of $\mathcal{H}_{s,s-1}(\bar{k},\alpha)$ has a unique monomial. Thus, the polynomials at $\bar{k}$ are linearly independent, so $\mathcal{H}_{s,s-1}$ is non-degenerate. Since every $\mathfrak{g}/\mathfrak{g}^{(i+1)}$ is step–$s$ or lower, the same argument will show that $\mathcal{H}_{s,i}$ is non degenerate for all $i<s$. We conclude that $\mathfrak{g}$ is ss–great. ∎
Combining Proposition 5.5 with Theorem 3.7 gives us Theorem A. We also show that $m$-greatness is preserved by taking products and factors.
@@
For clarity, in this proof we will let $\mathcal{H}^{\mathfrak{g}}_{m,p}$ be the map $\mathcal{H}_{m,p}$ associated to the Lie group $\mathfrak{g}.$ Similarly, let $\alpha_{ij}^{\mathfrak{g}}$ be the Malcev coordinates of the projection into $\mathfrak{g}$ of the $V_{i}$.
For the first part, observe that $\mathcal{H}_{m,p}^{\mathfrak{g}_{1}\times\mathfrak{g}_{2}}(k,\alpha^{\mathfrak{g}_{1}\times\mathfrak{g}_{2}})=(\mathcal{H}_{m,p}^{\mathfrak{g}_{1}}(k,\alpha^{\mathfrak{g}_{1}}),\mathcal{H}_{m,p}^{\mathfrak{g}_{2}}(k,\alpha^{\mathfrak{g}_{2}}))$. Since both of the $\mathfrak{g}_{i}$’s are $m$-great, $\mathcal{D}(\mathcal{H}^{\mathfrak{g}_{i}}_{m,p})$ is a positive codimension variety. Thus, there exists a $k_{0}$ in $\mathcal{D}(\mathcal{H}^{\mathfrak{g}_{1}}_{m,p})^{c}\cap\mathcal{D}(\mathcal{H}^{\mathfrak{g}_{2}}_{m,p})^{c}$. It follows that $k\in\mathcal{D}(\mathcal{H}_{m,p})$. Thus, $\mathcal{H}_{m,p}^{\mathfrak{g}_{1}\times\mathfrak{g}_{2}}$ is non-degenerate. Since $p$ was arbitrary, we conclude that $\mathfrak{g}_{1}\times\mathfrak{g}_{2}$ is $m$-great.